The partial pressure of gas X is 340 mm Hg. The partial pressure of gas Y is 264 mm Hg. The partial pressure of gas Z is 100 torr. What is the total pressure exerted by the gases in this system?


1
Expert's answer
2021-04-16T05:10:34-0400

According to Dalton's Law of Partial Pressure;

Pt= Px +Py+ Pz

= (340+264+100)mmHg (since 1torr=1mmHg)

= 704mmHg